# DisCO: Reinforcing Large Reasoning Models with Discriminative Constrained Optimization

### v2

@@ 5.1 A basic approach
Motivated by the connection with AUC maximization, we redesign the objective directly from the principle of discriminative learning. For a given question $q$, let $s_{\theta}(o,q)$ denote a scoring function that measures how likely the model $\pi_{\theta}$ “predicts” the output $o$ for a given input $q$ 111in the context of generative models, “predicts” is like “generates”. . Then the AUC score for the “task” $q$ is equivalent to $\mathbb{E}_{o\sim\pi_{\text{old}}^{+},o^{\prime}\sim\pi_{\text{old}}^{-}}[\mathbb{I}(s_{\theta}(o,q)>s_{\theta}(o^{\prime},q))]$. Using a continuous surrogate function $\ell$, we form the following objective (in expectation form) for maximization:
@@
Different surrogate functions $\ell(\cdot)$ can be used. For comparison, with GRPO, we simply use the identity function $\ell(s)=s$. One difference from the discriminative objective (6) is that we use a single scoring function $s_{\theta}(o,q)$ for both positive outputs $o$ and negative outputs $o^{\prime}$. It is notable that the different scoring functions for positive and negative outputs in (6) actually arise from the clipping operations of GRPO objective. Recent works have found that the clipping could lead to entropy collapse [79]. In addition, the clipping could cause the vanishing gradient, which may also slow down the learning process. To avoid these issues, we consider non-clipping scoring functions.
@@
Scoring functions. We consider two choices of scoring functions, i.e., log-likelihood and likelihood ratio. The log-likelihood (log-L) scoring function is defined by $s_{\theta}(o,q)=\frac{1}{|o|}\sum_{t=1}^{|o|}\log\pi_{\theta}(o_{t}|q,o_{<t})$. The likelihood ratio (L-ratio) scoring function is computed by $s_{\theta}(o,q)=\frac{1}{|o|}\sum_{t=1}^{|o|}\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\text{old}}(o_{t}|q,o_{<t})}$. In Appendix B.2, we discuss about the connection between the two scoring functions and the surrogate objectives of vanilla policy gradient methods [68] and TRPO [56], respectively.
@@
Stabilize the training with Constrained Optimization. Training instability is a long-standing issue in RL [56, 57]. Different methods have been introduced to ensure stability. Recent RL-based methods for learning reasoning models either follow the clipping operation of PPO [57] or use the KL divergence regularization $\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$ or $\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})$ [59, 61, 52]. However, the clipping operation could lead to the entropy collapse [79], which we try to avoid by using the non-clipped scoring function. The KL divergence regularization $\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$ while being used in traditional RL is not effective for preventing entropy collapse [26, 79]. Regarding the regularization with $\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})$, earlier studies [56, 57] has found that it would be difficult to choose a single value of the regularization parameter that performs well across different problems or even within a single problem where the the characteristics change over the course of learning. To tackle this issue, we revisit the idea of trust region constraint of TRPO [56], i.e., restricting the updated model $\theta$ in the trust region $\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})\leq\delta$. As a result, we solve the following discriminative constrained optimization problem:
$\max_{\theta}\mathcal{J}_{1}(\theta):=\mathbb{E}_{q}\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q),o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)}\ell(s_{\theta}(o,q)-s_{\theta}(o^{\prime},q))$ (8)
@@
For sake of efficiency, we use a different optimization approach from TRPO to solve the above constrained optimization. Inspired by the recent advances of non-convex ineqaulity constrained optimization algorithm [35], we adopt a squared-hinge penalty function for the constraint and solve the following problem with an appropriate penalty parameter $\beta$:
@@
where $[\cdot]_{+}=\max\{\cdot,0\}$. It has been shown that under an appropriate assumption regarding the constraint function and $\beta$, solving the above squared-hinge penalized objective (9) can return a KKT solution of the original constrained problem (8). We refer the readers to [35] for more in-depth analysis of this approach.
@@ 5.2 An improved approach for tackling imbalanced rollouts
One advantage of designing the objective based on the principle of discriminative learning is the ability to leverage a wide range of advanced techniques from the literature to improve training. A key challenge in RL fine-tuning for reasoning models is the sparse rewards, which leads to imbalance in generated rollouts. Specifically, for some questions where $p(q)\ll 1$, the number of negative outputs can significantly exceed the number of positive ones. This reflects a classic data imbalance issue, which has been extensively studied in the discriminative learning community [84, 58, 50]. To address this issue, we consider distributionally robust optimization (DRO) [84, 50].
Let us first discuss why the basic approach could be ineffective for combating the imbalanced rollouts. The objective function $\mathcal{J}_{1}$ is motivated by maximizing AUC for each question $q$, i.e., $\mathbb{E}_{o\sim\pi_{\text{old}}^{+},o^{\prime}\sim\pi_{\text{old}}^{-}}[\mathbb{I}(s_{\theta}(o,q)>s_{\theta}(o^{\prime},q))]$. However, when there is much more negative data than positive data, AUC is not a good measure. For example, let us consider a scenario that there are 1 positive $o_{+}$ and 100 negatives $\{o_{-}^{1},\ldots,o_{-}^{100}\}$. If the scores of these data are $s(o^{1}_{-},q)=0.9,s(o_{+},q)=0.5,s(o^{2}_{-},q)=s(o^{3}_{-},q)\ldots=s(o^{100}_{-},q)=0.001$, then the AUC score is $\frac{99}{100}=0.99$. The AUC score is high but is not informative as the model still generates the negative data $o^{1}_{-}$ more likely than the positive data $o_{+}$. In the literature, this issue has been addressed by maximizing a partial AUC score, which considers the pairwise order between all positives and the top ranked negatives. We utilize a surrogate function of partial AUC score formulated from the perspective of DRO [84].
Consider a question $q$ and a positive data $o$. We denote by $\mathcal{Q}$ the set of probability measures $Q$ on negative data given $q$ (absolutely continuous with respect to $\pi_{\text{old}}^{-}(\cdot|q)$). Denote by $\mathbb{D}_{\text{KL}}(\mathcal{Q},\pi_{\text{old}}^{-}(\cdot|q))$ the KL divergence between a distribution $Q$ and the negative data distribution $\pi_{\text{old}}^{-}(\cdot|q)$. A DRO formulation for partial AUC maximization is given by [84][Theorem 2]:
@@
It is easy to show that $\mathcal{J}_{2}(\theta)\leq\mathcal{J}_{1}(\theta)$ by Jensen’s inequality for the convex function $-\log$. Hence, maximizing $\mathcal{J}_{2}(\theta)$ will automatically increasing $\mathcal{J}_{1}(\theta)$. However, the reverse is not true. This also explains why maximizing $\mathcal{J}_{2}(\theta)$ could be more effective than maximizing $\mathcal{J}_{1}(\theta)$. The above risk function is also known as optimized certainty equivalents (OCE) in mathematical finance [4]. We would like to point out that although OCE or DRO has been considered for RL in existing works [65, 73], they differ from our work for addressing different issues. Wang et al. [65] apply the OCE to compute a robust reward, replacing the standard expected reward used in traditional RL settings. Xu et al. [73] adopt DRO for direct preference optimization of LLMs, aiming to mitigate the noise in human preference data by addressing the distributional shift between the empirical distribution of $(q,o,o^{\prime})$ and its true underlying distribution.
@@
GRPO Yes Yes regularization, $\pi_{\text{ref}}$ clipped L-ratio No
Dr. GRPO Yes Yes No clipped L-ratio No
DAPO Yes Yes No clipped L-ratio No
GPG Yes No No log-L No
TRPA No No regularization, $\pi_{\text{old}}$ log L-ratio No
@@
$\max_{\theta}\mathcal{J}_{2}(\theta):=-\mathbb{E}_{q}\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\tau\log\bigg{(}\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)}\exp\bigg{(}\frac{s_{\theta}(o^{\prime},q)-s_{\theta}(o,q)}{\tau}\bigg{)}\bigg{)},$ (11)
@@
To differentiate the approach for solving (8) and (11), we refer to the former as DisCO-b and the latter as DisCO. In practice, all expectations will be replaced by empirical averages and the KL divergence is also estimated at each iteration by using sampled data following [52]. We present a full algorithm in Algorithm 1 in Appendix. Finally, we give a comparison between DisCO and existing RL fine-tuning methods for reinforcing LRMs from different aspects in Table 1.
@@ 6 Experiments
In this section, we empirically evaluate the effectiveness of the proposed DisCO by comparing with GRPO and other variants for reinforcing an SFT-finetuned model.
Table 2: Comparison with baseline models and baseline methods for fine-tuning 1.5B models. OpenAI-o1-preview is included as a reference. MRL denotes Max Response Length utilized in training/testing. The shaded models are trained by other works and the shaded numbers are reported in their original works or in [42]. All other results are either evaluated on existing models or on the models trained by us using different approaches. Methods in the bottom area are all for fine-tuning DeepSeek-R1-Distill-Qwen-1.5B model on the same DeepScaleR dataset. DS is short for DeepSeek-R1, DSR is short for DeepScalaR.
@@
OpenAI-o1-Preview - 0.4 - 0.814 - - - -
DS-Distill-Qwen-1.5B 32k+ / 32k 0.288 0.263 0.828 0.629 0.265 0.433 0.451
DS-Distill-Qwen-1.5B 32k+ / 8k 0.181 0.215 0.758 0.515 0.237 0.353 0.376
STILL-3-1.5B-preview 29k / 32k 0.325 0.248 0.844 0.667 0.290 0.454 0.471
DSR-1.5B-Preview 24k / 32k 0.431 0.304 0.878 0.736 0.302 0.500 0.525
DSR-1.5B-Preview 24k / 8k 0.358 0.258 0.860 0.679 0.297 0.473 0.488
@@
Table 3: Comparison with baseline models and baseline methods for fine-tuning 7B models. Methods in the bottom area are all for fine-tuning DeepSeek-R1-Distill-Qwen-7B model on the the same DeepScalaR dataset.
@@
DS-Distill-Qwen-7B 32k+ / 32k 0.560 0.396 0.923 0.825 0.380 0.568 0.609
DS-Distill-Qwen-7B 32k+ / 8k 0.402 0.292 0.873 0.688 0.355 0.471 0.513
GRPO-LEAD-7B 8k / 8k 0.470 0.345 0.893 0.748 0.372 0.500 0.555
TRPA 8k / 8k 0 .570 - 0.870 0.780 0.360 0.550 -

### v4

@@ 5.1 A basic approach
Motivated by the connection with AUC maximization, we redesign the objective directly from the principle of discriminative learning. For a given question $q$, let $s_{\theta}(o,q)$ denote a scoring function that measures how likely the model $\pi_{\theta}$ “predicts” the output $o$ for a given input $q$ 222in the context of generative models, “predicts” is like “generates”. . Then the AUC score for the “task” $q$ is equivalent to $\mathbb{E}_{o\sim\pi_{\text{old}}^{+},o^{\prime}\sim\pi_{\text{old}}^{-}}[\mathbb{I}(s_{\theta}(o,q)>s_{\theta}(o^{\prime},q))]$. Using a continuous surrogate function $\ell$, we form the following objective (in expectation form) for maximization:
@@
Different surrogate functions $\ell(\cdot)$ can be used. For comparison, with GRPO, we simply use the identity function $\ell(s)=s$. One difference from the discriminative objective (6) is that we use a single scoring function $s_{\theta}(o,q)$ for both positive outputs $o$ and negative outputs $o^{\prime}$. It is notable that the different scoring functions for positive and negative outputs in (6) actually arise from the clipping operations of GRPO objective. Recent works have found that the clipping could lead to entropy collapse yu2025dapo. In addition, the clipping could cause the vanishing gradient, which may also slow down the learning process. To avoid these issues, we consider non-clipping scoring functions.
@@
Scoring functions. We consider two choices of scoring functions, i.e., log-likelihood and likelihood ratio. The log-likelihood (log-L) scoring function is defined by $s_{\theta}(o,q)=\frac{1}{|o|}\sum_{t=1}^{|o|}\log\pi_{\theta}(o_{t}|q,o_{<t})$. The likelihood ratio (L-ratio) scoring function is computed by $s_{\theta}(o,q)=\frac{1}{|o|}\sum_{t=1}^{|o|}\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\text{old}}(o_{t}|q,o_{<t})}$. In Appendix B.2, we discuss the connection between the two scoring functions and the surrogate objectives of vanilla policy gradient methods williams1992simple and TRPO schulman2015trust, respectively.
@@
Stabilize the training with Constrained Optimization. Training instability is a long-standing issue in RL schulman2015trust; schulman2017proximal. Different methods have been introduced to ensure stability. Recent RL-based methods for learning reasoning models either follow the clipping operation of PPO schulman2017proximal or use the KL divergence regularization $\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$ or $\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})$ shao2024deepseekmath; su2025trust; ouyang2022training. However, the clipping operation could lead to the entropy collapse yu2025dapo, which we try to avoid by using the non-clipped scoring function. The KL divergence regularization $\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})$ while being used in traditional RL is not effective for preventing entropy collapse skywork-or1-2025; yu2025dapo. Regarding the regularization with $\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})$, earlier studies schulman2015trust; schulman2017proximal has found that it would be difficult to choose a single value of the regularization parameter that performs well across different problems or even within a single problem where the the characteristics change over the course of learning. To tackle this issue, we revisit the idea of trust region constraint of TRPO schulman2015trust, i.e., restricting the updated model $\theta$ in the trust region $\mathbb{D}_{\text{KL}}(\pi_{\text{old}}||\pi_{\theta})\leq\delta$. As a result, we solve the following discriminative constrained optimization problem:
$\max_{\theta}\mathcal{J}_{1}(\theta)=\mathbb{E}_{q}\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q),o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)}\ell(s_{\theta}(o,q)-s_{\theta}(o^{\prime},q))$ (8)
@@
For sake of efficiency, we use a different optimization approach from TRPO to solve the above constrained optimization. Inspired by the recent advances of non-convex ineqaulity constrained optimization algorithm li2024model, we adopt a squared-hinge penalty function for the constraint and solve the following problem with an appropriate penalty parameter $\beta$:
@@
where $\cdot]_{+}=\max\{\cdot,0\}$. It has been shown that under an appropriate assumption regarding the constraint function and $\beta$, solving the above squared-hinge penalized objective (9) can return a KKT solution of the original constrained problem (8). We refer the readers to li2024model for more in-depth analysis of this approach.
@@ 5.2 An improved approach for tackling imbalanced rollouts
One advantage of designing the objective based on the principle of discriminative learning is the ability to leverage a wide range of advanced techniques from the literature to improve training. A key challenge in RL fine-tuning for reasoning models is the sparse rewards, which lead to imbalance in generated rollouts. Specifically, for some questions where $p(q)\ll 1$, the number of negative outputs can significantly exceed the number of positive ones. This reflects a classic data imbalance issue, which has been extensively studied in the discriminative learning community pmlr-v162-zhu22g; DBLP:journals/corr/Shalev-ShwartzW16; 10.5555/3294996.3295057. To address this issue, we consider distributionally robust optimization (DRO) pmlr-v162-zhu22g; 10.5555/3294996.3295057.
Let us first discuss why the basic approach could be ineffective for combating the imbalanced rollouts. The objective function $\mathcal{J}_{1}$ is motivated by maximizing AUC for each question $q$, i.e., $\mathbb{E}_{o\sim\pi_{\text{old}}^{+},o^{\prime}\sim\pi_{\text{old}}^{-}}[\mathbb{I}(s_{\theta}(o,q)>s_{\theta}(o^{\prime},q))]$. However, when there is much more negative data than positive data, AUC is not a good measure. For example, let us consider a scenario where there are 1 positive $o_{+}$ and 100 negatives $\{o_{-}^{1},\ldots,o_{-}^{100}\}$. If the scores of these data are $s(o^{1}_{-},q)=0.9,s(o_{+},q)=0.5,s(o^{2}_{-},q)=s(o^{3}_{-},q)\ldots=s(o^{100}_{-},q)=0.001$, then the AUC score is $\frac{99}{100}=0.99$. The AUC score is high but is not informative as the model still generates the negative data $o^{1}_{-}$ more likely than the positive data $o_{+}$. In the literature, this issue has been addressed by maximizing a partial AUC score, which considers the pairwise order between all positives and the top ranked negatives. We utilize a surrogate function of partial AUC score formulated from the perspective of DRO pmlr-v162-zhu22g.
Consider a question $q$ and a positive data $o$. We denote by $\mathcal{Q}$ the set of probability measures $Q$ on negative data given $q$ (absolutely continuous with respect to $\pi_{\text{old}}^{-}(\cdot|q)$). Denote by $\mathbb{D}_{\text{KL}}(Q,\pi_{\text{old}}^{-}(\cdot|q))$ the KL divergence between a distribution $Q$ and the negative data distribution $\pi_{\text{old}}^{-}(\cdot|q)$. A DRO formulation for partial AUC maximization is given by pmlr-v162-zhu22g[Theorem 2]:
@@
It is easy to show that $\mathcal{J}_{2}(\theta)\leq\mathcal{J}_{1}(\theta)$ by Jensen’s inequality for the convex function $\log$. Hence, maximizing $\mathcal{J}_{2}(\theta)$ will automatically increasing $\mathcal{J}_{1}(\theta)$. However, the reverse is not true. This also explains why maximizing $\mathcal{J}_{2}(\theta)$ could be more effective than maximizing $\mathcal{J}_{1}(\theta)$. The above risk function is also known as optimized certainty equivalents (OCE) in mathematical finance oce. We would like to point out that although OCE or DRO has been considered for RL in existing works wang2025reductionsapproachrisksensitivereinforcement; xu2025distributionallyrobustdirectpreference, they differ from our work for addressing different issues. Wang et al. wang2025reductionsapproachrisksensitivereinforcement apply the OCE to compute a robust reward, replacing the standard expected reward used in traditional RL settings. Xu et al. xu2025distributionallyrobustdirectpreference adopt DRO for direct preference optimization of LLMs, aiming to mitigate the noise in human preference data by addressing the distributional shift between the empirical distribution of $(q,o,o^{\prime})$ and its true underlying distribution.
@@
GRPO guo2025deepseek Yes Yes regularization, $\pi_{\text{ref}}$ clipped L-ratio No
Dr. GRPO liu2025understanding Yes Yes No clipped L-ratio No
DAPO yu2025dapo Yes Yes No clipped L-ratio No
GPG chu2025gpg Yes No No log-L No
TRPA su2025trust No No regularization, $\pi_{\text{old}}$ log L-ratio No
@@
$\max_{\theta}\mathcal{J}_{2}(\theta)=-\mathbb{E}_{q}\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q)}\tau\log\bigg(\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)}\exp\bigg(\frac{s_{\theta}(o^{\prime},q)-s_{\theta}(o,q)}{\tau}\bigg)\bigg),$ (11)
@@
To differentiate the approach for solving (8) and (11), we refer to the former as DisCO-b and the latter as DisCO. In practice, all expectations will be replaced by empirical averages and the KL divergence is also estimated at each iteration by using sampled data following ouyang2022training. We present a full algorithm in Algorithm 1 in Appendix. Finally, we give a comparison between DisCO and existing RL fine-tuning methods for reinforcing LRMs from different aspects in Table 1.
@@ 6 Experiments
In this section, we empirically evaluate the effectiveness of the proposed DisCO by comparing with GRPO and other variants for reinforcing SFT-finetuned models.
Table 2: Comparison with baseline models and baseline methods for fine-tuning 1.5B models. OpenAI-o1-preview is included as a reference. MRL denotes Max Response Length utilized in training/testing. The shaded models are trained by other works and the shaded numbers are reported in their original works or in deepscaler2025. All other results are either evaluated on existing models or on the models trained by us using different approaches. Methods in the bottom area are all for fine-tuning DeepSeek-R1-Distill-Qwen-1.5B model on the same DeepScaleR dataset. DS is short for DeepSeek-R1, DSR is short for DeepScaleR.
@@
[…]
@@
Table 3: Comparison with baseline models and baseline methods for fine-tuning 7B models. Methods in the bottom area are all for fine-tuning DeepSeek-R1-Distill-Qwen-7B model on the same DeepScaleR dataset.
@@
[…]
@@
Table 4: Comparison with baseline models and baseline methods for fine-tuning 8B models. Methods in the bottom area are all for fine-tuning DeepSeek-R1-Distill-Llama-8B model on the same DeepScaleR dataset.
[…]
GRPO 8k / 8k 0.410 0.240 0.873 0.759 0.307 0.506 0.516
GRPO+ER 8k / 8k 0.408 0.277 0.882 0.785 0.311 0.511 0.529
Dr. GRPO 8k / 8k 0.423 0.285 0.867 0.786 0.300 0.497 0.526
DAPO 8k / 8k 0.333 0.308 0.879 0.794 0.325 0.522 0.527
TRPA 8k / 8k 0.454 0.279 0.864 0.756 0.289 0.518 0.527
DisCO (L-ratio) 8k / 8k 0.506 0.356 0.900 0.831 0.326 0.553 0.579
DisCO (log-L) 8k / 8k 0.523 0.354 0.896 0.843 0.331 0.560 0.584
Task Setting. We validate our method on mathematical reasoning tasks. Specifically, we use the DeepScaleR-Preview-Dataset deepscaler2025 for training, which includes AIME problems from 1984 to 2023, AMC problems before 2023, and questions from the Omni-MATH gao2024omni and Still min2024imitate datasets, totaling approximately 40.3k unique problem-answer pairs. We evaluate models on six benchmark datasets: AIME 2024, AIME 2025, MATH 500 hendrycks2021measuring; lightman2023let, AMC 2023, Minerva lewkowycz2022solving, and Olympiad Bench (O-Bench) he2024olympiadbench. Following guo2025deepseek; deepscaler2025, we adopt the pass@1 metric chen2021evaluating averaged over $k=16$ responses for each question to ensure the reliability of model performances. The metric for each question is calculated as $\frac{1}{k}\sum_{i=1}^{k}\mathbb{I}(o_{i}\ \text{is correct})$, where $o_{i}$ denotes the $i$-th generated response. For both the training and evaluation of our method and the baselines (unless otherwise specified), the maximum response length is limited to 8k tokens. To verify the generalizability of our method to other datasets, we also conducted experiments on DAPO-Math-17k yu2025dapo dataset, which is included in the Appendix A.3.
Models. We conduct experiments with fine-tuning three models: DeepSeek-R1-Distill-Qwen-1.5B model (Q1.5B), DeepSeek-R1-Distill-Qwen-7B model (Q7B), and DeepSeek-R1-Distill-Llama-8B (L8B). All are distilled reasoning models.
Baselines. We primarily compare our methods with five most recent state-of-the-art reinforcement learning methods, including (1) GRPO guo2025deepseek; (2) GRPO with an entropy regularization (GRPO-ER) that adds an entropy on probabilities of output tokens as a regularization to prevent entropy collapse, which is used by DeepScaleR deepscaler2025; (3) Dr. GRPO liu2025understanding; (4) DAPO’s objective yu2025dapo; (5) TRPA su2025trust. For a comprehensive evaluation, we also include a set of reasoning models that are trained from the same base model by other studies with various techniques, such as (6) STILL-3-1.5B-preview chen2025empirical, which adapt GRPO by periodically replacing the reference model after a fixed number of training steps; (7) DeepScaleR(DSR)-1.5B-Preview that uses maximum response length of 24k for training deepscaler2025; (8) GRPO-LEAD-7B zhang2025grpo, which extends GRPO by incorporating length-dependent rewards, explicit penalty terms, and difficulty-based advantage reweighting to encourage concise and precise reasoning.
Training Details. For all the methods, we tune the constant learning rate in $5\mathrm{e}^{-7},1\mathrm{e}^{-6},2\mathrm{e}^{-6}]$ with AdamW optimizer with weight decay as 0.01. Generally, a learning rate of $2\text{e}^{-6}$ works better for the Q1.5B model, $1\text{e}^{-6}$ for the Q7B model, and $5\text{e}^{-7}$ for the L8B model. We employ a training batch size of 128, a mini-batch size of 32, and 8 responses for each question. The temperature is set to 0.6 for both training and evaluation, following the usage recommendation from guo2025deepseek. For GRPO, $\beta$ is set to 0.001 as commonly used chen2025empirical; deepscaler2025. For GRPO-ER, we use a coefficient of 0.001 for the entropy regularization deepscaler2025. For DAPO, we set $\epsilon_{low}$ to 0.2 and $\epsilon_{high}$ to 0.28 by following their paper. For our method, $\delta$ is set to $10^{-4}$ based on the empirical observation that the average KL divergence is around $2*10^{-5}$ and $\beta$ is set to $10^{3}$ such that the effective weight of the KL regularization when the constraint is violated by $\delta$ is on the order of $\beta*\delta=0.1$. Since L-ratio and log-L scoring functions have different orders, we choose $\tau=1$ for L-ratio and $\tau=10$ for log-L scoring function, from $\{0.5,1,5,10\}$. For fair comparisons, we do not implement Dynamic Sampling yu2025dapo for DAPO and other methods, as it introduces approximately three times the sampling cost at each training step. All methods are run for 1,400 steps on Q1.5B models and 1,000 steps on Q7B/L8B models. Evaluations are conducted every 200 steps, and the best performance for each method is reported.